ID |
Date |
Author |
Type |
Category |
Subject |
12123
|
Fri May 20 00:06:19 2016 |
Varun | Update | General | SURF 2016 |
I have written a basic version of AGC, and have done some tests with a data file. will do tests on whitening and agc today. Also, today I have to go to the SSN office. Hence will be late.
-Varun
Quote: |
Finished writing the code on whitening. I have to still test it. uploaded on github noise cancellation repo. @eric could you give me some data of noise power spectral density for testing the code?
-Varun
Quote: |
Hello, I am Varun Kelkar. I will be working at the 40m lab as a SURF student this summer with Eric Quintero on Audio processing for real time control system signals. This week I will mostly be working on implementing basic DSP C-code offline. Currently I am trying to write a code for noise whitening.
-Varun
|
|
|
12122
|
Thu May 19 16:29:20 2016 |
Steve | Update | endtable upgrade | Optical layout almost complete |
|
Attachment 1: ETMX_4x3_closed.jpg
|
|
Attachment 2: sealedETMXenclosure.jpg
|
|
12121
|
Wed May 18 17:42:52 2016 |
Varun | Update | General | SURF 2016 |
Finished writing the code on whitening. I have to still test it. uploaded on github noise cancellation repo. @eric could you give me some data of noise power spectral density for testing the code?
-Varun
Quote: |
Hello, I am Varun Kelkar. I will be working at the 40m lab as a SURF student this summer with Eric Quintero on Audio processing for real time control system signals. This week I will mostly be working on implementing basic DSP C-code offline. Currently I am trying to write a code for noise whitening.
-Varun
|
|
12120
|
Wed May 18 01:10:22 2016 |
gautam | Update | COC | Finesse modelling |
I've been working on putting together a Finesse model for the current 40m configuration. The idea was to see if I could reproduce a model that is in agreement with what we have been seeing during the recent DRFPMI locks. With Antonio and EricQs help, I've been making slow progress in my forays into Finesse and pyKat. Here is a summary of what I have so far.
- Arm lengths were taken from some recent measurements done by yutaro and me
- Recycling cavity lengths were taken from Gabriele's elog 9590 - it is likely that the lengths I used have errors ~1cm - more on this later. Furthermore, I've tried to incorporate the flipped RC folding mirrors - the point being to see if I can recover, for example, a power recycling gain of ~7 which is what was observed for the recent DRFPMI locks.
- I used Yutaro's most recent arm loss numbers, and distributed it equally between ITM and ETM for modeling purposes.
- For all other optics, I assumed a generic loss number of 25ppm for each surface
Having put together the .kat file (code attached, but this is probably useless, the new model with RC folding mirrors the right way will be what is relevant), I was able to recover a power recycling gain of ~7.5. The arm transmission at full lock also matches the expected value (125*80uW ~ 10mW) based on a recent measurement I did while putting the X endtable together. I also tuned the arm losses to see (qualitatively) that the power recycling gain tracked this curve by Yutaro. EricQ suggested I do a few more checks:
- Set PRM reflectivity to 0, scan ETMs and look at the transmission - attachment #1 suggests the linewidth is as we expect
- Set ETM reflectivity to 0, scan PRM - attachment #2 suggests a Finesse of ~60 for the PRC which sounds about right
- Set ETM reflectivity to 0, scan SRM and verify that only the 55 MHz sidebands resonate - Attachment #3
Conclusion: It doesn't look like I've done anything crazy. So unless anyone thinks there are any further checks I should do on this "toy" model, I will start putting together the "correct" model - using RC folding mirrors that are oriented the right way, and using the "ideal" RC cavity lengths as detailed on this wiki page. The plan of action then is
- Evaluating the mode-matching integrals between the PRC and the arm cavities as a function of the radius of curvature of PR2 and PR3
- Same as above for the SRC
- PRC gain as a function of RoC of folding mirrors
- Mode overlap between the modes from the two arm cavities as a function of the RoC of the two ETMs (actually I guess we can fix RoC of ETMy and just vary RoC of ETMx).
Sidenote to self: It would be nice to consolidate the most recent cavity length measurements in one place sometime... |
Attachment 1: arms.pdf
|
|
Attachment 2: PRC.pdf
|
|
Attachment 3: SRC.pdf
|
|
Attachment 4: Finesse_model.zip
|
12119
|
Tue May 17 14:46:51 2016 |
Steve | Update | VAC | RGA scan at day 595 |
We have good RGA scan now. There was no scan for 3 months. |
Attachment 1: RGAscan595d.png
|
|
Attachment 2: pd78-560Hz-d600.png
|
|
12118
|
Tue May 17 05:50:43 2016 |
Varun Kelkar | Update | General | SURF 2016 |
Hello, I am Varun Kelkar. I will be working at the 40m lab as a SURF student this summer with Eric Quintero on Audio processing for real time control system signals. This week I will mostly be working on implementing basic DSP C-code offline. Currently I am trying to write a code for noise whitening.
-Varun |
12117
|
Sun May 15 19:48:08 2016 |
Steve | Update | VAC | run out of N2 |
3-4 hrs ago we run out of nitrogen. We are back to Vacuum Normal
|
Attachment 1: noN2.png
|
|
12116
|
Thu May 12 14:29:58 2016 |
gautam | Update | VAC | RGA back up and running |
It looks like the hardware reset did the trick. Previously, I had just tried ssh-ing into c0rga and rebooting it. At the time, however, Steve and I noticed that the various LEDs on the RGA unit weren't on, as they are supposed to be in the nominal operating state. Today, Steve reported that all LEDs except the RS232 one were on today, so I just tried following the steps in this elog again, looks like things are back up and running. I'm attaching a plot of the scan generated using plotrgascan MATLAB script, it looks comparable to the plot in elog 11697, which if I remember right, was acceptable.
Unless there is some reason we want to keep this c0rga machine, I will recommission one of the spare Raspberry Pis lying around to interface with the RGA scanner when I get the time...
Quote: |
Quote: |
Our last RGA scan is from February 14, 2016 We had a power outage on the 15th
Gautom has not succeded reseting it. The old c0rga computer looks dead. Q may resurrect it, if he can?
|
The c0rga computer was off, I turned it on via front panel button. After running RGAset.py, RGAlogger.py seems to run. However, there are error messages in the output of the plotrgascan MATLAB script; evidiently there are some negative/bogus values in the output.
I'll look into it more tomorrow.
|
This is a cold scan. |
Attachment 1: RGAscan_12May2016.png
|
|
12115
|
Wed May 11 16:39:01 2016 |
ericq | Update | VAC | c0rga alive, output wonky |
Quote: |
Our last RGA scan is from February 14, 2016 We had a power outage on the 15th
Gautom has not succeded reseting it. The old c0rga computer looks dead. Q may resurrect it, if he can?
|
The c0rga computer was off, I turned it on via front panel button. After running RGAset.py, RGAlogger.py seems to run. However, there are error messages in the output of the plotrgascan MATLAB script; evidiently there are some negative/bogus values in the output.
I'll look into it more tomorrow. |
12114
|
Tue May 10 03:44:59 2016 |
ericq | Update | LSC | Relocked |
ALSX noise is solidly within past acceptable performance levels. The DRFPMI was locked on four out of six attempts.
Some housekeeping was done:
- PMC aligned
- Static alignment voltages of X end PZT mirrors offloaded by turning mount screws
- Rough comissioning of AUX X dither alignment
- Locking scripts reverted to AUX X Innolight voltage/temperature sign convention
The recombination of the QPD signals to common / differential is imperfect, and limited how well we could keep the interferometer aligned, since the QPD at X has changed. This needs some daytime work.
Some sensing matrix measurements were made, to be meditated upon for how to 1F the DRMI.
Other to-dos:
- Bandpass + notch combo for green refl PDs
- SRCL, and to a lesser extant, MICH feedforward subtraction (See DARM vs. other length DOF coherence plot below)
- Fiber couple AUX X light
- Make IFO work good

As an aside, Gautam and I noticed numerous green beams coming from inside the vacuum system onto the PSL table. They exist only when green is locked to the arms. Some of them come out at very non-level angles and shine in many places. This doesn't make me feel very happy; I suppose we've been living with it for some time. |
Attachment 1: 2016-05-10_DARMcoherence.pdf
|
|
12113
|
Sun May 8 08:39:21 2016 |
rana | Update | LSC | Green PDH demod lowpass |
Indeed. This is why the LSC PDs have a 2f notch in addition to the 1f resonance. In recent versions, we also put a 2f notch in the feedback of the preamp which comes after the diode but before the mixer. The overall 1f to 2f ratio that we get is in the 50-60 dB region. I don't think we have to go that far with this thing; having a double LC already seems like it should be pretty good, or we could have a single LC bandpass with a 2f notch all in one Pomona box. |
12112
|
Sat May 7 09:40:40 2016 |
ericq | Update | LSC | Green PDH demod lowpass |
As I was looking at filter designs, it seemed difficult to get 40dB of supression at 2F with a bandpass without going to a pretty high order, which would mean a fair number of lossy inductors.
I'll keep working on it. Maybe we don't need 40dB... |
12111
|
Fri May 6 19:08:52 2016 |
rana | Update | LSC | Green PDH demod lowpass |
Seems weird to design a PD lowpass with a corner at the modulation frequency. Recall what our strategy is with the other photodetectors we use for PDH servos: bandpass, not low-pass, and the band has to be wide enough to not effect the phase of the servo. |
12110
|
Fri May 6 16:42:12 2016 |
ericq | Update | LSC | Green PDH demod lowpass |
I've build the filter, and it seems to have the desired TF shape.

I also re-purposed the 70k lowass to a ~120k lowpass by changing the 68nF caps to 22nF caps, since we still want some post-mixer rolloff.
However, putting the ELPF in the chain caused some weird shapes in the OLG. I still need to get to the bottom of it. However, just with the post-mixer LPF modification, here's what the OLG looks like:

As Rana surmises, we definitely still add a boost and maintain a 10k UGF. I still need to look into the state of the remote boost.... |
12109
|
Thu May 5 21:28:44 2016 |
gautam | Update | endtable upgrade | Innolight PZT capacitance |
I suggested in an earlier elog that after the repair of the NPRO, the PZT capacitance may have changed dramatically. This seems unlikely - I measured the PZT capacitance with the BK Precision LCR meter and found it to be 2.62 nF, which is in excellent agreement with the numbers from elogs 3640 and 4354 - but this makes me wonder how the old setup ever worked. If the PZT capacitance were indeed that value, then for the Pomona box design in elog 4354, and assuming the PM at ~216kHz which was the old modulation frequency was ~30rad/V as suggested by the data in this elog, we would have had a modulation depth of 0.75 if the Function Generator were set to output a Signal at 2Vpp (2Vpp * 0.5 * 0.05 * 30rad/V = 1.5rad pp)! Am I missing something here?
Instead of using an attenuator, we could instead change the capacitor in the pomona box from 47pF mica to 5pF mica to realize a modulation depth of ~0.2 at the new modulation frequency of 231.25 kHz. In any case, as elog 4354 suggests, the phase introduced by this high-pass filter is non-zero at the modulation frequency, so we may also want to install an all-pass filter which will allow us to control the demodulation phase. This should be easy enough to implement with an Op27 and passive components we have in hand...
|
12108
|
Thu May 5 14:05:01 2016 |
rana | Update | endtable upgrade | ALS status update |
All seems very fishy. Its not good to put attenuators and filters in nilly-willy.
- Once the post-PD bandpass has been designed and constructed, you should be able to use whatever PD gain setting gives you the best SNR. There's no need to use more PD gain than necessary; it just reduces the PD bandwidth. What is the input referred current noise of the PD at the different gain settings?
- The open loop mixer output *should* be very large. It should be reduced to mV only when the loop is closed.
- The better way to estimate the modulation depth is to lock the arm on red as usual and then scan the EX laser and look at the green transmission. The FSR is 3.7 MHz, so the SBs should show up well in a narrow scan around the carrier.
- I guess its going to be tough to impedance match the splitter box to the NPRO PZT, since its impedance is all over the place at 200-300 kHz, but you could put a 50 Ohm in-line terminator in there somewhere?
- The Bode plot seems to indicate that we could easily get a 10 kHz UGF and then switch on a Boost. Is the remote Boost switch disabled or always ON? I am suspicious of the plot and think that the coarse trace is probably missing some sharp resonances which will sneakily bite you.
|
12107
|
Thu May 5 14:03:52 2016 |
ericq | Update | LSC | Further Aux X PDH tweaks |
This morning I poked around with the green layout a bit. I found that the iris immediately preceding the viewport was clipping the ingoing green beam too much, opening it up allowed for better coupling to the arm. I also tweaked the positions of the mode matching lenses and did some alignment, and have since been able to achieve GTRX values of around 0.5.
I also removed the 20db attenuator after the mixer, and turned the servo gain way down and was able to lock easily. I then adjusted the gain while measuring the CLG, and set it where the maximum gain peaking was 6dB, which worked out to be a UGF of around 8kHz. On the input monitor, the PDH horn-to-horn voltage going into the VGA is 2.44V, which shouldn't saturate the G=4 preamp stage of the AD8336, which seems ok.
The ALS sensitivity is now approaching the good nominal state:

There remains some things to be done, including comprehensive dumping of all beams at the end table (especially the reflections off of the viewport) and the new filters to replace the current post-mixer LPF, but things look pretty good. |
Attachment 1: 2016-05-05_newals.pdf
|
|
12106
|
Thu May 5 04:05:03 2016 |
ericq | Update | LSC | Aux X PDH checks |
We took an OLG measurement of the green PDH loop. It seems consistent with past measurements. I've added a trace for the the post-mixer lowpass, to show its contribution to the phase loss. (EDIT: updated with measured LPF TF)

I used this measured OLG and the datasheet laser PZT conversion factor to calibrate the control signal monitor into the AUX laser frequency noise, it looks consistent with the frequency noise measured via the PSL PLL (300 Hz/rtHz @ 100Hz). Above a few tens of kHz, the control signal measurement is all analyzer noise floor, due to the fourth order 70kHz lowpass after the mixer (the peaks change height significantly depending on the analyzer input range, so I don't think they're on the laser). Gautam will follow up with more detailed measurements of both the error and control signals as he noisebudgets, this was just intended as a quick consistency check.

|
12105
|
Thu May 5 03:05:37 2016 |
gautam | Update | endtable upgrade | ALS status update |
[ericQ, gautam]
Today we spent some time looking into the PDH situation at the X end. A summary of our findings.
- There is something that I don't understand with regards to the modulation signal being sent to the laser PZT via the sum+HPF pomona box - it used to be that with 2Vpp signal from the function generator, we got ~5mVpp signal at the PZT, which with the old specs resulted in a modulation of ~0.12rad. Now, however, I found that there was a need to place a 20dB attenuator after the splitter from the function generator in order to realize a modulation depth of ~0.25 (which is what we aim for, measured by locking to the TEM00 modes of the carrier and sidebands and comparing the ratio of powers). It could be that the PZT capacitance has changed dramatically after the repair. Nevertheless, I still cant reconcile the numbers. We measured the transfer function from the LO input of the pomona box to the output with the PZT connected, and figure there should be ~70dB of attentuation (with the 20dB additional attenuator in place). But this means 1Vpp*0.0003*70rad/V = 0.02rad which is an order of magnitude away from what the ratio of powers suggest. Maybe the measurement technique was not valid. In any case, this setup appears to work, and I'm also able to send +7dBm to the mixer which is what it wants (function generator output is 3Vpp).
- In addition to the above, I found that the demodulated error signal had a peak-to-peak of a few volts. But the PDH servo is designed to have tens of mV at the input. Hence, it was necessary to turn down the gain of the REFL PD to 10dB and add a 20dB attenuator between mixer output and servo input.
- While Johannes and I were investigating this earlier in the afternoon, we found that the waveform going to the laser PZT was weirdly distorted (still kind of sinusoidal in shape, but more rounded, I will put up a picture shortly). This may not be the biggest problem, but perhaps there is a better way to pipe the LO signal to the PZT and mixer than what is currently done.
- We then looked at loop transfer function and spectrum of the control signal. Plots to follow. They look okay.
- I measured the green power coming onto the PSL table. It is ~400uW. After optimizing alignment, the green transmission is ~0.4 according to whatever old normalization we are using.
- We then recovered the X green beatnote and looked at the ALS noise spectrum. Beatnote amplitude at the beat PD is ~ -27dBm. The coherence in the region of a few hundred Hz suggests that some improvements can be made to the PDH situation (the gain of the PDH servo is maxed out at the X end at the moment...). But the bottom line is this is probably good enough to get back to locking...
|
Attachment 1: ALS_noiseSpec_5May2016_2.pdf
|
|
Attachment 2: Coherence_5May2016.pdf
|
|
Attachment 3: image.jpeg
|
|
12104
|
Mon May 2 19:14:18 2016 |
gautam | Update | endtable upgrade | Optical layout almost complete |
With Steve's help, I installed the Oplev earlier today. I adjusted the positions of the two lenses until I deemed the spot size on the QPD satisfactory by eye. As a quick check, I verified using the DTT template that the UGF is ~5Hz for both pitch and yaw. There is ~300uW of power incident on the QPD (out of ~2mW from the HeNe). In terms of ADC counts, this is ~13,000 counts which is about what we had prior to taking the endtable apart. There are a couple of spots from reflections off the black glass plate in the vacuum chamber, but in general, I think the overall setup is acceptable.
This completes the bulk of the optical layout. The only bits remaining are to couple the IR into the fiber and to install a power monitoring PD. Pictures to follow shortly.
Now that the layout is complete, it remains to optimize various things. My immediate plan is to do the following:
- Maximize green transmission by tweaking alignment. I should also do a quick check using mirror specs to see that the measured transmitted green power compares favourably to what is expected.
- Check the green PDH loop transfer function at the X end - this will allow me to set the gain on the uPDH box systematically.
- Re-establish green beats, check noise performance.
- There are possibly multiple beam dumps that have to be installed. For now, I've made sure that no high power IR beams are incident on the enclosure. But there are a couple of red and green beams that have to be accounted for.
I will also need to upload the layout drawing to reflect the layout finally implemented.
Not directly related:
The ETMx oplev servo is now on. I then wanted to see if I could lock both arms to IR. I've managed to do this successfully - BUT I think there is something wrong with the X arm dither alignment servo. By manually tweaking the alignment sliders on the IFOalign MEDM screen, I can get the IR transmission up to ~0.95. But when I run the dither, it drives the transmission back down to ~0.6, where it plateaus. I will need to investigate further.
GV Edit: There was some confusion while aligning the Oplev input beam as to how the wedge of the ETM is oriented. We believe the wedge is horizontal, but its orientation (i.e. thicker side on the right or left?) was still ambiguous. I've made a roughly-to-scale sketch (attachment #1) of what I think is the correct orientation - which turns out to be in the opposite sense of the schematic pinned up in the office area.. Does this make sense? Is there some schematic/drawing where the wedge orientation is explicitly indicated? My search of the elog/wiki did not yield any.. |
Attachment 1: ETMX_wedge.pdf
|
|
12103
|
Mon May 2 17:11:55 2016 |
rana | Update | COC | RC folding mirrors |
Antonio/Gautam are now developing a more up to date Finesse model of our recycling cavities to see what we can have there before our power recycling gain or cavity geometric stability is compromised. Expect that we will here a progress report on the model on Wednesday.
Some thoughts:
- RC folding mirrors need to be dichroic to allow green beams to get out.
- We should look at the specs Jamie used to get the RC folding mirrors last time and figure out what went wrong / what specs to change.
- T_1064 < 100 ppm. Hopefully < 50 ppm.
- On the AR side, we mainly want low AR for green, but nothing special for 1064, since that's taken care of by the HR.
- How much should we wedge these things?
- Should the wedge be horizontal?
- Can we get someone in Downs to update the optical layout?
- What microroughness do we need?
- The mirrors must be flat, with the 500 m < RoC < 100 km. Part of the Finesse modeling is to figure out what happens if the RoC is in the 300 - 1000 m range. Better stability?
|
12102
|
Mon May 2 17:06:58 2016 |
rana | Summary | COC | G&H optics to Fullerton/HWS for anneal testing |
Steve sent 4 of our 1" diameter G&H HR mirrors to Josh Smith at Fullerton for scatter testing. Attached photo is our total stock before sending. |
Attachment 1: 20160427_182305.jpg
|
|
12101
|
Fri Apr 29 16:13:36 2016 |
ericq | Update | LSC | Green PDH demod lowpass |
We can get as much, if not more, attenuation of the 1F line in the mixer output that we get from the post-mixer LPF from using the following passive filter between the PD and mixer RF input:

There should still be some kind of LPF after the mixer, but I haven't yet determined what it should be; this will determine how much phase the PDH loop wins. At most, this should win around 25 degrees at 10kHz.
The filter was designed by referencing the "Handbook of Filter Synthesis" by Zverev, looking for an elliptic filter for matched source and load impedences, 40dB min attenuation in the stopband, a stopband frequency that starts at twice the corner frequency, and minimizing the VSWR between the PD and filter in the passband.
In terms of the tables in the book, this means: n=5, rho=2%, theta=30deg, K**2 = 1.0. The dimensionless component values were scaled by the corner frequency of 200kHz, and reference impedence of 50 Ohm. (The corner is a little lower than the real modulation frequency, since the nonzero resistance of the inductors pushes the frequency up a bit)
The ideal capactior values do not correspond to things we have in hand, so I checked our stock and chose the closest value to each one.Unsurprisingly, due to these component substitutions, and the fact that the coilcraft inductors have a resistance of about 7 Ohms, the predicted TF of the realizable filter does not match the design filter exactly. However, the predicition still looks like it will meet the requirement of 40dB of supression of the 2F line in the PD signal. (Since we have tunable inductors, I've used the ideal inductor values in generating the TF. In practice I'll inspect the TF while I tune them)
|
Desired |
Realizable |
C1 |
8.28 nF |
10 nF |
C2 |
1.39 nF |
1.5 nF |
C3 |
19.6 nF |
22 nF |
C4 |
4.22 nF |
4.7 nF |
C5 |
6.08 nF |
6.8 nF |
L2 |
43.1 nH |
32-48 nH + 7 Ohm |
L4 |
34.4 nH |
32-48 nH + 7 Ohm |

[In this TF plot, I've multiplied the real response by 2 to account for the voltage division that occurs with ideal 50 Ohm impedance matching, to make 0dB the reference for proper matching]
The filter's phase delay at the modulation frequency is just about 180, which as a time delay of 5usec works out to 9 degrees of phase loss at 10kHz in the PDH loop. According to some old measurements, the current LPF costs something like 35 degrees at 10k, so this wins at most around 25 degrees, depedent on what LPF we put after the mixer.
LISO source both traces is attached! |
Attachment 3: elp_liso.zip
|
12100
|
Fri Apr 29 16:05:23 2016 |
gautam | Update | endtable upgrade | Cleaning ETMX vacuum dirty window |
After a second round of F.C. application, I think the window is clean enough and there are no residual F.C. pieces anywhere near the central parts of the window (indeed I think we got most of it off). So I am going to go ahead and install the Oplev.
Quote: |
It looks very promising.
|
|
Attachment 1: IMG_0755.JPG
|
|
12099
|
Fri Apr 29 00:55:46 2016 |
gautam | Update | endtable upgrade | green PDH locked to Xarm |
Using the modulation frequency suggested here, I hooked up the PDH setup at the X-end and succeeded in locking the green to the X arm. I then rotated the HWP after the green Faraday to maximize TRX output, which after a cursory alignment optimization is ~0.2 (I believe we were used to seeing ~0.3 before the end laser went wonky). Obviously much optimization/characterization remains to be done. But for tonight, I am closing the PSL and EX laser shutters and applying first contact to the window once more courtesy more PEEK from Koji's lab in W Bridge. Once this is taken care of, I can install the Oplev tomorrow, and then set about optimizing various things in a systematic way.. MC autolocker has also been disabled...
Side note: for the IR Transmon QPD, we'd like a post that is ~0.75" taller given the difference in beam height from the arm cavity and on the endtable. I will put together a drawing for Steve tomorrow.. |
12098
|
Thu Apr 28 18:53:05 2016 |
rana | Update | LSC | Green PDH demod lowpass |
OK - but give us a circuit diagram and the expected before/after loop plots. Got to make sure we keep the right impedance from PD to mixer. Some of the Thorlabs PDs have a 50 Ohm instead of 0 Ohm source impedance. Maybe you can try it out now since the green arm is ready. |
12097
|
Thu Apr 28 15:23:11 2016 |
ericq | Update | LSC | Green PDH demod lowpass |
The 2F product out of the mixer is a natural concern when demodulating. However, I think this isn't so big of a deal in our green PDH servos; 420kHz isn't so high of a frequency that the servo amplifiers are bandwidth or slew-rate limited. Furthermore, the amplitude of this line is supressed by the loop somewhat, since it arises from the same field product that the loop is acting on. Measuring the Y end mixer output with a high impedance probe and the AG4395 shows it to be something like -50dBm.
In fact, the main thing that the pomona LPFs are accomplishing right now is filtering the 1F content of the mixer output that arises from the second order sideband creating a signal at 2F, and beating with the LO at (2F-1F)=1F. This line is something like -30dBm (5mVrms) at the mixer output; I can reproduce this amplitude with a back-of-the envelope calculation using a modulation depth of 0.3, 8V out of the PD at DC when unlocked, the mixer datasheet, and the nominal cavity parameters.
The nice thing about this is that we don't need to filter this after the mixer, we can use a [bandpass/lowpass/notch] filter before the mixer (as is done in the LSC demod boards) to filter out the 2F (420kHz) content of the PD signal, which will only introduce some small amount of linear time delay to the PDH loop, instead of the wicked phase loss from the current post-mixer LPF. We can then replace that 70kHz filter with something of lower order or higher corner frequency to win a good deal of phase in the PDH loop. |
12096
|
Thu Apr 28 08:49:47 2016 |
Steve | Update | PEM | possible noise sources schedule |
Building: Campus (see attached Map)
Date: Manhole 1 - May 3 through May 5
Manhole 2 – May 6 through May 10
Manhole 2 - May 16 through May 19
Manhole 3 – May 11 through May 19
Time: Noise: 7:00 a.m. To 5:00 p.m.
Access: 24 Hours a day
Interruption: Noise/Vehicular & Pedestrian Access
Storm Drain Manholes
*In order to repair 3 manholes associated with a large storm drain that runs north-south through the campus, work will take place at the
3 manholes shown on the map. This work will interrupt vehicular and pedestrian access on the paths adjacent to the manholes. Though the work at Manholes 1 and 2 will allow vehicular and pedestrian access around the manholes, the work at Manhole 3 will completely block the driveway running south from the southeast corner of Parking Lot 11. Noise will also be created by the repair
|
Attachment 1: Campus_B&W_Map-2.pdf
|
|
12095
|
Thu Apr 28 00:41:08 2016 |
gautam | Update | endtable upgrade | more progress - Transmon PD installed |
The IR Transmon system is almost completely laid out, only the QPD remains to be installed. Some notes:
- The "problem" with excessive green power reflected from the harmonic separator has been resolved. It is just very sensitive to the angle of incidence. In the present configuration, there is ~10uW of green power reflected from either side, which shouldn't be too worrisome. But this light needs to be dumped. Given the tiny amount, I think a black glass + sticky tape solution is best suited, given the space constraints. This does not reach the Transmon PDs because there is a filter in the path that is transmissive to IR only.
- I aligned the transmitted beam onto the Thorlabs PD, and reconnected the signal BNC cable (the existing cable wasn't long enough so I had to use a barrel connector and a short extension cable). I then reverted the LSC trigger for the X arm back to TRX DC and also recompiled c1ass to revert to TRX for the dither alignment. At the moment, both arms are stably locked, although the X arm transmission is saturated at ~0.7 after running the dither alignment. I'm not sure if this is just a normalization issue given the new beam path or if there is something else going on. Further investigations tomorrow.
- It remains to dump some of the unwanted green light from the addition of the harmonic separator...
- We may want to redesign some (or all) of the Transmon path - the lens currently in use seems to have been chosen arbitrarily. Moreover, it is quite stubbornly dirty, there are some markings which persist after repeated first contact cleaning...
I feel like once the above are resolved, the next step would be to PDH lock the green to the arm and see what sort of transmission we get on the PSL table. It may be the polarization or just alignment, but for some reason, the transmitted green light from the X arm is showing up at GTRY now (up to 0.5, which is the level we are used to when the Y arm has green locked!). So a rough plan of action:
- Install transmon QPD
- PDH lock green to X arm
- Fix the window situation - as Steve mentioned in an earlier elog, the F.C. cleaning seems to have worked well, but a little remains stuck on the window (though away from where any laser radiation is incident). This is resolved easily enough if we apply one more layer of F.C., but the bottle-neck right now is we are out of PEEK which is what we use to remove the F.C. once dried. Steve thinks a fresh stock should be here in the next couple of days...
- Once 3 is resolved, we can go ahead and install the Oplev.
- Which leaves the lst subsystem, coupling to the fiber and a power monitor for the NPRO. I have resolved to do both these using the 1% transmitted beam after the beamsplitter immediately after the NPRO rather than pick off at the harmonic separator after the doubling oven. I need to do the mode-matching calculation for coupling into the fiber and also adjust the collimating lens...
- Clean-up: make sure cables are tied down, strain-relieved and hooked up to whatever they are supposed to be hooked up to...
|
12094
|
Wed Apr 27 15:04:47 2016 |
Steve | Update | endtable upgrade | Cleaning ETMX vacuum dirty window |
It looks very promising.
|
Attachment 1: 1cETMX-Tcmp.jpg
|
|
12093
|
Wed Apr 27 14:06:31 2016 |
rana | Summary | General | meeting notes |
- Gautam will get help from Johannes and finish EX table by Monday.
- Steve will spend a day this week with Johannes on Green Monster bakeout.
- Q to analyzed green PDH servo and design demod low pass. Should we use the double LC notches to notch the 2f product? What's the demod filter attenuation requirement?
- Koji will make a drawing of the ruby suspension standoff prism and post into the elog so that Steve can get some quotes next week.
- Rana to implement 40m configuration in FOGprime17 and analyze RoC matching of ETMs. Get Antonio's help to analyzed SRC stability. Maybe use PyKat and Finesse since Antonio knows that stuff.
- Give OCXO boxes to WB refcav people. Rana get Rich to make another couple of boxes for 40m PMC, FSS, IMC.
- Rana/Koji get EKG to make specs and procure some new folding mirrors for the PRC/SRC. Make them a bit concave and dichroic.
|
12092
|
Wed Apr 27 09:45:56 2016 |
rana | Update | SUS | spare SOS tower |
Bah, we need ruby slippers for all future suspensions. Prism with curved backside and smooth grooves.
No aluminum, no cry.
Quote: |
Earth quake stops need viton tips.
Wirestandoffs are still aluminum.
|
|
12091
|
Wed Apr 27 09:05:10 2016 |
Steve | Update | General | AP viewport |
Sad situation
The anti-symmetric port
spider webs fly in the wind |
Attachment 1: APg.jpg
|
|
Attachment 2: AP.jpg
|
|
Attachment 3: APspiderWebs.jpg
|
|
12090
|
Tue Apr 26 23:19:42 2016 |
gautam | Update | endtable upgrade | Green aligned to arm - high order mode flashes seen |
Attachment #1
Layout as of today. Most of the green path is done. The Green REFL PD + PZT mirrors have not been hooked up to their respective power sources yet (I wonder if it's okay to start laying cables through the feedthroughs on either end of the table already, or if we want to put whatever it is that makes it airtight eventually in first?). A rough power budget has been included (with no harmonic separator just before the window), though some optimization can be done once the table is completely repopulated.
Attachment #2
A zoomed-in version of the REFL path.
Some general notes:
- I've tried to use the custom 3/4" O.D. posts + baseplate arrangement wherever possible (only 1 steering mirror is on a 1" post clamped with a fork to the table because of space constraints). Where the baseplates could not be bolted onto the table directly, I've used Newport SS Dogs to do the job.
- I checked for continuity between the PZT outer case and the table top with a multimeter, and found none. So I chose to leave the Thorlabs baseplates in place. For the REFL PD, I've used an insulating baseplate given to me by Steve.
- I've used some custom length 3/4" O.D. posts to get the beam up to the right height (~4.75") just before sending the green beam in. The beam height is 4" elsewhere.
- I was playing around with positioning the harmonic separator immediately before the vacuum chamber window - I found that there is a substantial amount of green light that is reflected, though there doesn't seem to be any IR leaking through. The mirror was labelled Y1-1037-45P, which is a code for CVI mirrors, though I believe it is a LaserOptik product and that we have a couple of other such mirrors in the optics cabinet - though they are all 1". This document suggests that from the back side, there should be <0.1% reflection of green while on the front side it should be < 3%. I will have to hunt a little more for the specs, and measure the powers to see if they match the previously quoted numbers. In any case, I'll have to think of how to separate the (unwanted) reflected green and the transmitted IR from the cavity in the IR transmon path.
- There are some minor changes to the planned layout posted here - I will update these in due course once the Transmon path and Oplev have been set up.
I am closing the PSL shutter and the EX laser shutters for the night as I have applied a layer of first contact to the window for cleaning purposes, and we don't want any laser light incident on it. It may be that the window is so dirty that we may need multiple F.C. cleaning rounds, we will see how the window looks tomorrow...
|
Attachment 1: IMG_2219.JPG
|
|
Attachment 2: IMG_2220.JPG
|
|
12089
|
Tue Apr 26 15:22:35 2016 |
Steve | Update | endtable upgrade | Cleaning ETMX vacuum dirty window |
Gautom is progressing with the layout nicely. The X-arm transmission window have not seen cleaning for decades. This should be the time to do it. Here is picture of dirtiness.
It is not that simple... How much effort should we put in it? The hole table with 1W inno laser plus... set up now about ~500 lbs We can pull it off carefully, but it is not risk free.
We should look at our other signal port windows! Gautom's long reach able him to do the first contact cleaning without moving anything. It is great!
|
Attachment 1: ETMX-Tvp.jpg
|
|
Attachment 2: ETMX-TvpDetail.jpg
|
|
12088
|
Mon Apr 25 11:07:06 2016 |
Steve | Update | SUS | spare SOS tower |
Earth quake stops need viton tips.
Wirestandoffs are still aluminum. |
Attachment 1: ETMXreplacment.jpg
|
|
12087
|
Fri Apr 22 13:58:13 2016 |
Steve | Update | PEM | leaky roof is fixed |
Dan sealed the leak today.
|
Attachment 1: leakyRoof_(2).jpg
|
|
12086
|
Thu Apr 21 15:12:38 2016 |
Steve | Update | VAC | RGA is not working |
Quote: |
Steve pointed out that in the aftermath of the Nitrogen running out a couple of times last week, the RGA had shut itself off thinking that there was a leak and so it was not performing the scheduled scans once a day. So the data files from the scheduled scans were empty in the /opt/rtcds/caltech/c1/scripts/RGA/logs directory. The wiki page for getting it up and running again is up-to-date, but the script RGAset.py did not exist on the c0rga machine, which the RGA is communicating with via serial port. I copied over the script RGAset.py from rossa to c0rga and ran the script on that machine - but the error flags it returned were not all 0 (indicating some error according to the manual) - so I edited the script to send just the initialize command ('IN0') and commented out the other commands, after which I got error flags which were all 0. After this, I ran a manual scan using 'RGAlogger.py', and it appears that the RGA is now able to take scans again - I'm attaching a plot of the scan results. We've saved this scan as a reference to compare against after a few days.
|
Our last RGA scan is from February 14, 2016 We had a power outage on the 15th
Gautom has not succeded reseting it. The old c0rga computer looks dead. Q may resurrect it, if he can? |
12085
|
Thu Apr 21 14:25:52 2016 |
gautam | Update | endtable upgrade | Green light recovered |
I've made progress on the new layout up to the doubling oven. After doing the coarse alignment with the diode current to the NPRO at ~1A, I turned it back up to the nominal 2A. I then rotated the HWP before the IR Faraday such that only ~470mW of IR power is going into the doubler (the rest is being dumped on razor beam dumps). After tuning the alignment of the IR into the doubling oven using the steering mirror + 4 axis translation stage on which the doubling oven is mounted, I get ~3.2mW of green after the harmonic separator and a HR mirror for green. The mode looks pretty good to the eye (see attachment #1), and the conversion efficiency is ~1.45%/W - which is somewhat less than the expected 2%/W but in the ballpark. It may be that some fine tweaking of the alignment + polarization while monitoring the green power can improve the situation a little bit (I think it may go up to ~4mW, which would be pretty close to 2%/W conversion efficiency). The harmonic separator also seems to be reflecting quite a bit of green light along with IR (see attachment #2) - so I'm not sure how much of a correction that introduces to the conversion efficiency.
While doing the alignment, I noticed that some amount of IR light is actually transmitted through the HR mirrors. With ~500mW of incident light at ~45 degrees, this transmitted light amounts to ~2mW. Turns out that this is also polarization dependant (see attachment #3) - for S polarized light, as at the first two steering mirrors after the NPRO, there is no transmitted light, while for P-polarized light, which is what we want for the doubling crystal, the amount transmitted is ~0.5%. The point is, I think the measured levels are consistent with the CVI datasheet. We just have to take care find all these stray beams and dump them.
I will try and optimize the amount of green power we can get out of the doubler a little more (but anyway 3mW should still be plenty for ALS). Once I'm happy with that, I will proceed with laying out the optics for mode-matching the green to the arm. |
Attachment 1: IMG_6567.JPG
|
|
Attachment 2: IMG_6568.JPG
|
|
Attachment 3: CVI_reflectivity.jpeg
|
|
12084
|
Wed Apr 20 14:36:45 2016 |
Steve | Update | PEM | dusty lab |
It is worth wiping table top covers. Use isopropanol soaked lint free wipes.
|
Attachment 1: AP_accm.jpg
|
|
Attachment 2: wiped_AP.jpg
|
|
Attachment 3: temp-partic_6months.png
|
|
Attachment 4: 3_years_PEM.png
|
|
12083
|
Tue Apr 19 18:37:29 2016 |
gautam | Update | endtable upgrade | Laser swap + optical layout |
Summary of work done over the last two days
- Lightwave NPRO + controller moved to PSL table
- The interlock is not connected to the controller
- Controller is not powered
- Innolight NPRO + controller installed at endtable
- Interlock has been connected
- For initial alignment purposes, I'm running it at an injection current of 1.000A (~50mW of IR out of the NPRO)
- Temperature of crystal set to 31.66 degrees in anticipation of operation in the nominal state
- Laying out optics
- Given that the mode out of the NPRO is different from that from the Lightwave, the mode-matching had to be re-done
- Attachment #1 shows the mode-matching solution being implemented
- Current state - I've placed all the optics up to and including the doubling crystal + oven. Alignment through IR Faraday is pretty good, QWP+HWP angles optimized to maximize transmission through the Faraday (<10% loss). Oven has been hooked up to temperature controller, and is currently set to 36.3 degrees. Coarse alignment into doubling crystal done at lower power. Even with the low IR power, I am able to see some green. It remains to turn the injection current up and do the fine alignment + lens position tweaking to maximize the green power from the doubling crystal - with ~1W of power, assuming 2%/W SHG efficiency, we should be seeing 20 mW of green (which is probably way too much)
Immediate next steps:
- Some optimization to be done with regards to beam dumps for rejected beam from IR Faraday. Also double check to make sure that the reflected beam from L1 doesn't go back directly to the laser (at the moment it doesn't, is there a standard way to do this? I was trying to have the lens as close to normal incidence as possible, but I may not have been entirely successful which is why the reflected beam is not going straight back at the moment).
- Optimize mode-matching into the doubling crystal
- Once the desired green mode is obtained, continue with the rest of the layout
- Update CAD drawing to reflect new layout
|
Attachment 1: IR_modematch_19April2016_2.pdf
|
|
12082
|
Tue Apr 19 10:58:53 2016 |
Steve | Update | safety | safety scans at PSL & tables |
The PSL had one 1064 nm beam to be blocked around the north east side. The end enclosures are fine.
|
12081
|
Mon Apr 18 00:29:00 2016 |
gautam | Update | General | Beam profiling + injection current scan |
Summary
I've finished up the remaining characterization of the repaired 1W Innolight NPRO - the beamscan yielded results that are consistent with an earlier beam-profiling and also the numbers in the datasheet. The output power vs diode current plot is mainly for diagnostic purposes in the future - so the plot itself doesn't signify anything, but I'm uploading the data here for future reference. The methodology and analysis framework for the beamscan is the same as was used here.
Attachment #1 - Beam-scan results for X-direction
Attachment #2 - Beam-scan results for Y-direction
Attachment #3 - Beam profile using fitted beam radii
Attachment #4 - Beam-scan data
Attachment #5 - Output power vs Injection current plot
Even though I remember operating at a diode current of 2.1A at some point in the past, while doing this scan, attempting to increase the current above 2.07A resulted in the "Clamp" LED on the front turning on. According to the manual, this means that the internal current limiting circuitry has kicked in. But I don't think this is a problem as we don't really even need 1W of output power. This is probably an indicator of the health of the diode as well?
Attachment #6 - Output power vs Injection current data
It remains to redo the mode-matching into the doubling oven and make slight modifications to the layout to accommodate the new laser + beam profile.
I plan to do these in the morning tomorrow, and unless there are any objections, I will begin installing the repaired 1W Innolight Mephisto on the X endtable tomorrow (18 April 2016) afternoon.
|
Attachment 1: BeamScan_x.pdf
|
|
Attachment 2: BeamScan_y.pdf
|
|
Attachment 3: ZScan.pdf
|
|
Attachment 4: BeamScan.mat
|
Attachment 5: Innolight_Current_Scan.pdf
|
|
Attachment 6: Innolight_Current_Scan.mat
|
12080
|
Fri Apr 15 23:11:49 2016 |
gautam | Update | General | Innolight 1W moved to SP table |
I have moved the 1W Innolight + controller from the PSL table to the SP table for beam profiling. |
12079
|
Fri Apr 15 18:38:12 2016 |
gautam | Update | endtable upgrade | Lightwave health check - NO IMPROVEMENT |
I re-measured the power levels today.
We have ~205mW out of the NPRO, and ~190mW after the Faraday. It doesn't look like the situation is going to improve dramatically. I'm going to work on a revised layout with the Innolight as soon as I've profiled the beam from it, and hopefully, by Monday, we can decide that we are going ahead with using the Innolight. |
12078
|
Fri Apr 15 18:35:57 2016 |
gautam | Update | General | New AUX laser measurements |
I've performed the temperature sweep of PSL vs Innolight 1W AUX laser.
- I followed the procedure in this elog - started by turning of FSS and FSS Slow servos, closed the PSL shutter, noted down the value of PSL temperature
- As noted in elog 3759, there are multiple temperatures at which a beat can be found. I recorded all that I could find. The IR beat frequency was < 20MHz at the temperatures recorded (and had an amplitude of a few dBm, but I used a 20dB coupler to look at the signal on the HP spectrum analyzer
- The PMC unlocked each time I changed the PSL temperature, but the PMC autolocker worked for me every time
- We should use curve 3 in attachment 1, it is the most reliable set of temperatures at which a beat can be found
- PSL diode current was 2.100A, AUX laser diode current was 2.001A
- Attachment 2 is the data
It remains to measure the output power vs diode current, and the beam profile. I will do the latter on the SP table where there is a little more space. Because we have 1W from this NPRO, the knife-edge method requires a power meter that has a large dynamic range and is sensitive enough to profile the beam accurately. After consulting the datasheets of the power meters we have available (Scientech, Ophir and Coherent) together with Koji, I have concluded that the Coherent calorimeter will be suitable. Its datasheet claims it can accurately measure incident powers of up to 100uW, although I think the threshold is more like 5-10mW, but this should still be plenty to get sufficient resolution for a Gaussian intensity profile with peak intensity of 1W. We also checked that the maximum likely power density we are likely to have during the waist measurement process (1W in a beam of diameter 160um) is within the 6kW/cm^2 quoted on the datasheet. |
Attachment 1: PSL_AUX_TEMP_SCAN.pdf
|
|
Attachment 2: PSL_AUX_TEMP_SCAN.mat
|
12077
|
Fri Apr 15 03:02:44 2016 |
ericq | Update | General | New AUX laser measurements |
The free running PSL+AUX beat frequency noise spectrum has been measured via PLL. AUX laser PZT PM and AM responses were measured too.
 
Rough notes about these measurements:
Laser -> QWP -> HWP -> PBS -> 10% BS -> Beat
3.4Vpp out of PD, (40% contrast)
20dB Coupler, output to analyzer, coupled output to Mixer (-a few dBm, didn't check specifically)
Mixer: ZP-3+, BLP-5.1 at output
LO: OCXO @ 36MHz 13dBm->5dB Att-> +8dBm LO at Mixer
Got ~65mVpp out of Mixer
Mixer out -> SR560, LP 3Hz, G=500 -> Pomona Summing node -> Laser PZT
~30kHz UGF ~30 deg phase
Spectra, OLG via SR785 taken with free running PSL, anthropomorphic temperature servo. Data sheet calibration used for PZT. SR560 output noise dominates over analyzer, mixer, PD. Spectrum looks ok, I think.
PM measured with AG4395. High impedance probe used for laser PZT, otherwise couldn't lock. PM calibrated via mixer voltage span for fringe-to-fringe.
PSL beam blocked, AUX power increased to read 8.0V, AM measured with AG4395.
AM/PM doesn't look to dissimilar to old measurements on wiki. ~230kHz looks like a fine modulation freq.
Still to be done to AUX laser:
- joint PSL/AUX temperature sweeps
- Output power vs. diode current
- Beam profile |
Attachment 1: aux_innolight_freqPSD.pdf
|
|
Attachment 2: aux_innolight_AMPM.pdf
|
|
Attachment 3: aux_innolight_data.zip
|
12076
|
Thu Apr 14 17:30:18 2016 |
ericq | Update | General | PLL measurement ongoing |
Just a heads up that some equipment is hooked up at the PSL table for the repaired AUX laser PLL measurement, I plan to continue with it tonight.
I've taken a few spectra that, along with the PZT coefficient from the repair sheet, that suggest the noise level is ok (incoherent sum of AUX and PSL at about ~3e4 / f Hz/rtHz), but calibrated plots, etc. will follow in time. |
12075
|
Wed Apr 13 18:25:07 2016 |
gautam | Update | endtable upgrade | Lightwave health check |
[Koji,gautam]
Lightwave NPRO information:
Model: 126-1064-700
Serial Number: 337
Manufactured: December 1998!!
Details of checks performed:
Koji tuned the parameters on the laser controller and we observed the following:
- Turning "ADJ" to +10 and the pumping current all the way up to the maximum (2.62A) allowed us to recover an output power of 300mW, at a laser crystal temperature of ~45degrees
- The output power increased almost monotonically as a function of the laser crystal temperature - why? We were able to see powers as high as 250mW (at ADJ=0) for the maximum crystal temperature of ~60 degrees.
- We checked that we could believe the readout of the power meter by measuring the power using the Scientech power meter - we saw ~270mW after the Faraday with this meter, accounting for ~10% loss through the Faraday, this corresponds to an output power of 300mW (all this was done at ADJ=+10, DC=2.62A). I suspect that the display is dodgy though, because changing the Diode Current from 2.52A to 2.62A increased the output power by almost 100mW, which seems hard to believe?
- The Lightwave NPRO does not have heat dissipation fins attached - could this be affecting the power output somehow? In any case, this has to be rectified. So if we decide to keep the Lightwave NPRO, the layout will still need minor changes to accommodate the heat fins. Steve, do we have these in hand?
Way forward
Ericq has begun the characterization of the repaired Innolight. We checked that it outputs 1W of power. We will now have to perform the following measurements:
- Frequency noise using PLL
- AM/PM response of the PZT
- Laser power output as a function of diode current - this will be useful for diagnostic purposes in the future
- AUX temperature vs PSL temperature at which beatnotes can be found
- Waist measurement - the mode matching and optical layout upstream of the doubling oven at least will have to be modified significantly
All of these will have to be done before installing this laser at the endtable.
I believe the consensus as of now is to go ahead with carrying out the above measurements. Meanwhile, we will keep the Lightwave NPRO on and see if there is some miraculous improvement. So the decision as to whether to use the Innolight is deferred for a day or two. |
12074
|
Wed Apr 13 11:00:28 2016 |
Steve | Update | endtable upgrade | X endtable repopulation |
ETMX optical table is grounded to ETMX chamber through 1 Mohms
The doubling oven temp controller is installed to reach its cable.
|